# Imperial College London, London, SW7 2AZ, UK

### v1

@@ 1 Introduction
In this note, rather than proposing a new general mechanism to solve this difficult problem, we focus on trying to make progress by computing the entanglement entropy in a standard cosmological set-up and consider its time evolution.  [^1] For other approaches to entanglement entropy in cosmology see for instance Wall:2010jtc ; Maldacena:2012xp ; Brahma:2020zpk ; Boutivas:2023mfg ; DuasoPueyo:2024usw . It is well known that in a quantum mechanical system/environment set-up, an entanglement entropy can be computed by tracing out the contribution from the environment, breaking unitarity and allowing the possibility to have entropy increase (Zurek:2003zz, ).
@@
$S_{{\rm vN}}=-{\rm tr}\rho\ln\rho=-\sum_{i}p_{i}\ln p_{i},$ (1)
where the second equality obtains in a basis in which $\rho$ is diagonalized. If the state is pure ($\rho^{2}=\rho$), this entropy is zero. A non-zero entropy obtains by coarse graining, i.e. averaging over a bundle of microscopic variables, or by considering a subsystem of a closed system and tracing over all observables in its complement (the “environment”). We will be discussing the latter situation, in other words the physics of an open system coupled to an environment.
The entropy of the open system is then defined by,
$\rho_{{\rm red}}\left(\alpha\right)={\rm tr}_{\beta}\rho\left(\alpha,\beta\right),$ (2)
@@
In our cosmological application, where we assume that the fluctuations for the CMB modes were produced during inflation, the division between system and environment can be naturally thought of as provided by the comoving Hubble scale: modes with $k\leq\Lambda(t)$ are “system” modes, while those with $k>\Lambda(t)$ are “environment” modes, with $\Lambda(t)=a(t)H$, where $a(t)$ is the scale factor and and $H$ the Hubble parameter Burgess:2014eoa ; Burgess:2020tbq ; Brahma:2020zpk ; Burgess:2022nwu . We show that the entanglement entropy of the system modes grows monotonically during accelerated expansion ($\ddot{a}>0$) for both Gaussian and weakly non-Gaussian states. [^2] These results can be generalised to other theories provided that $\ddot{a}>0$, see Colas:2024xjy .
@@ 2 Decoherence and entropy increase
In quantum mechanics, decoherence refers to the loss of quantum coherence of a system due to the interaction with the environment (for a review see for instance (Zurek:2003zz, )). In quantum field theory, the phenomenon of decoherence is closely related to what is now called open effective field theories (EFTs). Open EFTs differ from standard EFTs in the sense that the states that are integrated out are not necessarily of higher energies but depend on the separation between the system and the environment (see (Burgess:2020tbq, ; Burgess:2022rdo, ) for recent reviews).
@@ 2.1 Decoherence in discrete bipartite systems
$|\psi\left(0\right)\rangle$ $=\left(c_{1}|\,+\,\rangle+c_{2}|-\rangle\right)|e_{0}\rangle,\qquad\quad\,|c_{1}|^{2}+\,|c_{2}|^{2}=1,$ (6)
@@
$\lambda=\lambda_{\pm}=\frac{1}{2}\left\{1\pm\left(1-4|c_{1}|^{2}|c_{2}|^{2}\left(1-|r(t)|^{2}\right)\right)^{1/2}\right\},$ (12)
@@ 2.2 Entropy increase
In any case, since the reduced density matrix has positive (non-zero) entropy at any time $t>0$ on average we must have $\langle dS/dt\rangle>0$ which means that $\langle d|r|/dt\rangle<0$.
@@ 3.1 Soft vs hard modes and the diffusion equation
${\rho}_{\Lambda}=\int\prod_{k>\Lambda}d\phi_{k}\langle\phi_{k}|\Psi\rangle\langle\Psi|\phi_{k}\rangle.$ (16)
@@
$\int\prod_{k\leq\Lambda}d\phi_{k}P_{\Lambda}\left(\phi\right)=1,\qquad\,P_{\Lambda}\left(\phi\right)=\int\prod_{k>\Lambda}d\phi_{k}P\left(\phi\right).$ (17)
@@
$\frac{\partial P_{\Lambda}\left(\phi\right)}{\partial\ln\Lambda}=V_{{\rm space}}\frac{H^{2}}{8\pi^{2}}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{k^{3}}\frac{d\Omega_{\Lambda}}{d\ln\Lambda}\frac{\partial^{2}P_{\Lambda}}{\partial\phi_{k}\partial\phi_{k}^{*}},$ (19)
@@
$\frac{\dot{\Lambda}}{\Lambda}=\frac{1}{H}\left(H^{2}+\dot{H}\right)=\frac{\ddot{a}}{aH}\qquad{\rm then}\qquad\ddot{a}\geq 0\implies\dot{S}\geq 0.$ (21)
@@ 3.2 Purity and Entropy of a Gaussian state
with $\phi_{k}$ being the Fourier components $\phi_{k}=\int\frac{d^{3}x}{\left(2\pi\right)^{3/2}}\phi\left(x\right)e^{-i{\bf k}\cdot{\bf x}}$ of the field (the reality of $\phi$ implies $\phi_{-{\bf k}}=\phi_{{\bf k}}^{*}$). We’ve also assumed above that the IR modes are uncorrelated initially. Also following the literature (such as (Burgess:2014eoa, ; Burgess:2022nwu, ) the latter of which we follow closely in this discussion), we take the dominant interaction to be linear in the IR field. Under this assumption, each momentum mode evolves independently. Also let us define the real Fourier component fields $\phi_{k}\equiv\left(\phi_{k}^{R}+i\phi_{k}^{I}\right)$, and the covariance matrix,
$\bm{\Sigma}={\rm tr}{\rho}\begin{bmatrix}\phi^{2}&\left\{\phi,\pi\right\}\\ \left\{\phi,\pi\right\}&\pi^{2}\end{bmatrix}$,
@@
$|\Psi(t_{0})\rangle=U(t_{0},-\infty)|\Omega\rangle,$ (22)
@@
$\Psi(\phi,t_{0})\equiv\langle\phi|\Psi(t_{0})\rangle=\int_{\rm BD}^{\varphi(t_{0})=\phi}d[\varphi]e^{iS[\varphi;t_{0}]}.$ (23)
@@
$\rho[\phi_{+},\phi_{-};t_{0}]\equiv\langle\phi_{+}|\rho(t_{0})|\phi_{-}\rangle=\Psi[\phi^{a};t_{0}]\Psi^{*}[\phi_{b};t_{0}],$ (24)
where the subindex $\pm$ specifies in which part of the contour the field is defined. The reduced density matrix between field eigenstates is obtained by integrating UV degrees of freedom. When the coupling between the short and long modes is linear we obtain at leading order a Gaussian density matrix,   [^8] A detailed derivation of how a Gaussian state emerges through coarse-graining of short-wavelength modes, using the path integral formalism, is presented in Appendix A.
$\langle\phi^{+}_{kR}|{\rho}_{{\rm red}}|\phi^{-}_{kR}\rangle=\sqrt{\frac{1}{2\pi\Sigma_{k11}^{R}}}\exp\left(-\frac{a_{k}}{2}\left(\phi_{kR}^{+}\right)^{2}-\frac{a_{k}^{*}}{2}\left(\phi^{-}_{kR}\right)^{2}+c_{k}\phi^{-}_{kR}\phi^{+}_{kR}\right)\times\left(R\rightarrow I\right),$ (25)
@@
$\gamma_{k_{\rm CMB}}^{{\rm end}}\simeq 2.1\times 10^{-35}\left(\frac{M_{{\rm P}}}{\rho_{{\rm inf}}^{1/4}}\right)^{4},$
which gives an entropy per CMB mode of $O(10)$ from the above formula.
Let us now consider how far these exact calculations (albeit in a Gaussian state) can be replaced by simply considering the diagonal element
$P\left(\phi\right):=\prod_{k}\langle\phi_{k}|{\rho}_{{\rm red}}|\phi{}_{k}\rangle=\prod_{k}P\left(\phi_{k}\right),\qquad P\left(\phi_{k}\right):=\sqrt{\frac{1}{2\pi\Sigma_{k11}}}\exp\left(-\frac{\phi_{k}^{2}}{2\Sigma_{11}}\right),$ (29)
@@
$\langle\phi|{\rho}_{{\rm red}}|\phi^{\prime}\rangle=\frac{1}{\sqrt{2\pi\Sigma_{11}}}\exp\left[-\frac{1}{2\Sigma_{11}}\phi_{s}^{2}-\frac{\det\Sigma}{2\Sigma_{11}}\phi_{{\rm a}}^{2}+i\frac{\Sigma_{12}}{\Sigma_{11}}\phi_{{\rm s}}\phi_{{\rm a}}\right].$ (30)
@@ 3.3 The second law in cosmology for a Gaussian state
$\frac{\partial}{\partial\eta}\det\mathbf{\Sigma}=2{\mathrm{Re}}\,\left[{\cal F}_{k}(\eta,\eta_{0})\right]\Sigma_{11},$ (31)
and (see eqn. (3.37) and the sentence below it in (Burgess:2022nwu, ) ), as pointed out in this reference ${\mathrm{Re}}\,\left[{\cal F}_{k}(\eta,\eta_{0})\right]$is expected to be positive on general grounds.  [^10] This of course would have followed from the properties of the Lindblad equation, in particular the necessity of preserving the trace and the positivity of the reduced density matrix. However here we do not have strict Markovianity so this is not guaranteed. For a detailed discussion of this including a generalization of the arguments of (Burgess:2022nwu, ), see (Lopez:2025arw, ). Explicitly in the model that is investigated in detail there,
@@
$\frac{\partial}{\partial\eta}\det\mathbf{\Sigma}>0,\qquad{\rm for}\qquad-k\eta\ll 1.$ (32)
So it seems that in this set up $\det\mathbf{\Sigma}$ increases without bound for late times and hence the purity appears to go to zero. However as pointed out in the paragraph after eqn. (28) in the previous subsection, purity at the end of inflation for a CMB mode is small but non-zero.
@@ 3.4 Second law in cosmology beyond the Gaussian approximation
$C_{B}^{*}\left(\eta,\eta^{\prime};{\bf x-x^{\prime}}\right)=C_{B}\left(\eta^{\prime},\eta;{\bf x^{\prime}-x}\right).$ (37)
@@
$\frac{\partial\rho}{\partial\eta}\simeq-\frac{1}{2}\int d^{3}x\int d^{3}x^{\prime}G^{2}(\eta)\left(\left\{v\left(\eta,{\bf x}\right)v\left(\eta,{\bf x}^{\prime}\right),\rho(\eta)\right\}-2v\left(\eta,{\bf x}\right)\rho(\eta)v\left(\eta,{\bf x}^{\prime}\right)\right)C_{B}\left({\bf x-x^{\prime}}\right).$ (39)
This has the form of a time-dependent Lindblad equation. From here we can compute the variation of the quantum entropy $S=-{\rm tr}\,\rho\,{\rm ln}\rho$ given by $\frac{dS}{d\eta}=-{\rm tr}\,\frac{d\rho}{d\eta}\ln\rho$.
Now let us assume that the background is spatially compact so that the density matrix can be expanded in a eigenbasis labeled by $\alpha,\beta$ etc. For instance in the absence of interaction (assumed to be the case in the far past) we have $\rho=\prod_{k_{\alpha}>\Lambda}\int_{\phi_{k}}\langle\phi_{k_{\alpha}}|\left\{\phi_{k^{\prime}}\right\}\rangle\langle\{\phi_{k^{\prime}}\}|\phi_{k_{\alpha}}\rangle$, where the $k_{\alpha}$’s take discrete set of values taking space to be compact as before. With time evolution this basis will of course get transformed, but given that the density matrix for the system is still Hermitian there will still be some basis in which $\rho$ will be diagonalized. To simplify the notation we label this basis $\left\{|\alpha\rangle\right\}$ and take it to be discrete with $\alpha|\beta>=\delta_{\alpha\beta}$. So $\rho=\sum_{\alpha}p_{\alpha}|\alpha\rangle\langle\alpha|,$ $\sum_{\alpha}p_{\alpha}=1,\,0<p_{\alpha}<1$.
@@
$\frac{dS}{d\eta}$ $\simeq\int d^{3}x\int d^{3}x^{\prime}G^{2}\left(\eta\right)\left(\sum_{\alpha\beta}(v(\eta,{\bf x})v(\eta,{\bf x}^{\prime}))_{\alpha\beta}\left(p_{\alpha}\ln\rho_{\alpha}\right)\delta_{\alpha\beta}+p_{\alpha}\left((v(\eta,{\bf x})v(\eta,{\bf x}^{\prime})\right)_{\alpha\beta}\left(\ln p_{\alpha}\right)\delta_{\beta\alpha}\right$.
@@
$=\sigma_{\beta\alpha}.$ (40)
@@
$\frac{dS}{d\eta}$ $\simeq G^{2}\sum_{\alpha\beta}\sigma_{\alpha\beta}\left(p_{\alpha}-p_{\beta}\right)\ln p_{\alpha}=\sum_{\alpha\beta}\sigma_{\beta\alpha}\left(p_{\beta}-p_{\alpha}\right)\ln p_{\beta}$
@@
In the second line above we used (40) and in the last line the positive semi-definiteness of $\sigma$, (41). We recall that the quantity $\left(p_{\alpha}-p_{\beta}\right)\left(\ln p_{\alpha}-\ln p_{\beta}\right)$ is positive as can be easily seen by using ${\rm ln}\,y\leq y-1$ for $y>1$. We then conclude that entropy increases monotonically.

### v2

@@ 1 Introduction
In this note, rather than proposing a new general mechanism to solve this difficult problem, we focus on trying to make progress by computing the entanglement entropy in a standard cosmological set-up and consider its time evolution  [^1] For other approaches to entanglement entropy in cosmology see for instance Wall:2010jtc ; Maldacena:2012xp ; Brahma:2020zpk ; Boutivas:2023mfg ; DuasoPueyo:2024usw .. It is well known that in a quantum mechanical system/environment set-up, an entanglement entropy can be computed by tracing out the contribution from the environment, breaking unitarity and allowing the possibility to have entropy increase (Zurek:2003zz, ).
@@
$S_{{\rm vN}}=-{\rm tr}\rho\ln\rho=-\sum_{i}p_{i}\ln p_{i}$ (1)
where the second equality obtains in a basis in which $\rho$ is diagonalized. If the state is pure ($\rho^{2}=\rho$), this entropy is zero. A non-zero entropy obtains by coarse graining, i.e. averaging over a bundle of microscopic variables, or by considering a subsystem of a closed system and tracing over all observables in its complement (the “environment”). We will be discussing the latter situation, in other words the physics of an open system coupled to an environment. The entropy of the open system is then defined in terms of,
$\rho_{{\rm red}}\left(\alpha\right)={\rm tr}_{\beta}\rho\left(\alpha,\beta\right)$ (2)
@@
In our cosmological application, where we assume that the fluctuations for the CMB modes were produced during inflation, the division between system and environment can be naturally thought of as provided by the comoving Hubble scale: modes with $k\leq\Lambda(t)$ are “system” modes, while those with $k>\Lambda(t)$ are “environment” modes, with $\Lambda(t)=a(t)H$ Lombardo:2005iz ; Campo:2008ju ; Campo:2008ij ; Burgess:2014eoa ; Burgess:2020tbq ; Brahma:2020zpk ; Burgess:2022nwu . We show that the entanglement entropy of the system modes grows monotonically during accelerated expansion ($\ddot{a}>0$) for both Gaussian and weakly non-Gaussian states. [^2] These results can be generalised to other theories provided that $\ddot{a}>0$, see Colas:2024xjy .
@@ 2 Decoherence and entropy increase
In quantum mechanics, decoherence refers to the loss of quantum coherence of a system due to the interaction with the environment (for a review see for instance (Zurek:2003zz, )). In quantum field theory, the phenomenon of decoherence is closely related to what is now called open effective field theories (EFTs). Open EFTs differ from standard EFTs in the sense that the states that are integrated out are not necessarily of higher energies but depend on the separation between the system and the environment (see (Burgess:2020tbq, ) for a review).
@@ 2.1 Decoherence in discrete bipartite systems
$|\psi\left(0\right)\rangle$ $=\left(c_{1}|\,+\,\rangle+c_{2}|-\rangle\right)|e_{0}\rangle,\qquad\quad\,|c_{1}|^{2}+\,|c_{2}|^{2}=1$ (6)
@@
$\lambda=\lambda_{\pm}=\frac{1}{2}\left\{1\pm\left(1-4|c_{1}|^{2}|c_{2}|^{2}\left(1-|r(t)|^{2}\right)\right)^{1/2}\right\}$ (12)
@@ 3.1 Soft vs hard modes and the diffusion equation
${\rho}_{\Lambda}=\int\prod_{k>\Lambda}d\phi_{k}\langle\phi_{k}|\Psi\rangle\langle\Psi|\phi_{k}\rangle$ (16)
@@
$\int\prod_{k\leq\Lambda}d\phi_{k}P_{\Lambda}\left(\phi\right)=1,\qquad\,P_{\Lambda}\left(\phi\right)=\int\prod_{k>\Lambda}d\phi_{k}P\left(\phi\right)$ (17)
@@
$\frac{\partial P_{\Lambda}\left(\phi\right)}{\partial\ln\Lambda}=V_{{\rm space}}\frac{H^{2}}{8\pi^{2}}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{k^{3}}\frac{d\Omega_{\Lambda}}{d\ln\Lambda}\frac{\partial^{2}P_{\Lambda}}{\partial\phi_{k}\partial\phi_{k}^{*}}$ (19)
@@
$\frac{\dot{\Lambda}}{\Lambda}=\frac{1}{H}\left(H^{2}+\dot{H}\right)=\frac{\ddot{a}}{aH}\qquad{\rm then}\qquad\ddot{a}\geq 0\implies\dot{S}\geq 0$ (21)
@@ 3.2 Purity and Entropy of a Gaussian state
with $\phi_{k}$ being the Fourier components $\phi_{k}=\int\frac{d^{3}x}{\left(2\pi\right)^{3/2}}\phi\left(x\right)e^{-i{\bf k}\cdot{\bf x}}$ of the field (the reality of $\phi$ implies $\phi_{-{\bf k}}=\phi_{{\bf k}}^{*}$). We’ve also assumed above that the IR modes are uncorrelated initially. Also following the literature (such as (Burgess:2014eoa, ; Burgess:2022nwu, ) the latter of which we follow closely in this discussion), we take the dominant interaction to be linear in the IR field. Under this assumption, each momentum mode evolves independently. Also let us define the real Fourier component fields $\phi_{k}\equiv\left(\phi_{k}^{R}+i\phi_{k}^{I}\right)$, and the matrix,
$\bm{\Sigma}={\rm tr}{\rho}\begin{bmatrix}\phi^{2}&\left\{\phi,\pi\right\}\\ \left\{\phi,\pi\right\}&\pi^{2}\end{bmatrix}$
@@
$|\Psi(t_{0})\rangle=U(t_{0},-\infty)|\Omega\rangle$ (22)
@@
$\Psi(\phi,t_{0})\equiv\langle\phi|\Psi(t_{0})\rangle=\int_{\rm BD}^{\varphi(t_{0})=\phi}d[\varphi]e^{iS[\varphi;t_{0}]}$ (23)
@@
$\rho[\phi_{+},\phi_{-};t_{0}]\equiv\langle\phi_{+}|\rho(t_{0})|\phi_{-}\rangle=\Psi[\phi^{a};t_{0}]\Psi^{*}[\phi_{b};t_{0}]$ (24)
where the subindex $\pm$ specifies in which part of the contour the field is defined. The reduced density matrix between field eigenstates is obtained by integrating UV degrees of freedom. When the coupling between the short and long modes is linear we obtain at leading order a Gaussian density matrix  [^8] A detailed derivation of how a Gaussian state emerges through coarse-graining of short-wavelength modes, using the path integral formalism, is presented in Appendix A.,
$\langle\phi^{+}_{kR}|{\rho}_{{\rm red}}|\phi^{-}_{kR}\rangle=\sqrt{\frac{1}{2\pi\Sigma_{k11}^{R}}}\exp\left(-\frac{a_{k}}{2}\left(\phi_{kR}^{+}\right)^{2}-\frac{a_{k}^{*}}{2}\left(\phi^{-}_{kR}\right)^{2}+c_{k}\phi^{-}_{kR}\phi^{+}_{kR}\right)\times\left(R\rightarrow I\right)$ (25)
@@
$\gamma_{k_{\rm CMB}}^{{\rm end}}\simeq 2.1\times 10^{-35}\left(\frac{M_{{\rm P}}}{\rho_{{\rm inf}}^{1/4}}\right)^{4}$
which gives an entropy per CMB mode of $O(10)$ from the above formula. In any case the period of inflation is finite and in general the FRW phase takes over long before $\eta$ goes to zero [^9] Note on duration of inflation: Assume that $H$ is constant during inflation (approximately true). We have, integrating the definition $H$ and taking (with no loss of generality) the end of inflation to be $t=0$ (where $t$ is proper time), $a(t)=a_{e}e^{Ht},\,t\leq 0$, $a_{e}$ being the scale factor at the end of inflation. As we mentioned in the introduction we are working in a cut off theory, and in any case the inflationary phase is not expected to extend all the way to the cosmological singularity (i.e. $a=0)$. So let us take the onset of inflation to be at scale factor $a_{i}$. Integrating the definition of conformal time $d\eta/dt=1/a(t)$ we have, $\eta-\eta_{i}=-1/\left(a(t)H\right)+1/\left(a_{i}H\right)$. Taking $\eta_{i}=-1/a_{i}H$ we have the usual expression $\eta=-1/\left(a(t)H\right)$. The duration of inflation is in any case finite, $\eta_{e}-\eta_{i}=-1/\left(a_{e}H\right)+1/\left(a_{i}H\right)$. So since during inflation $a$ is neither zero nor infinite our calculations have neither a UV nor an IR divergence..
Let us now consider how far these calculations (albeit in a Gaussian state) can be replaced by simply considering the diagonal element
$P\left(\phi\right):=\prod_{k}\langle\phi_{k}|{\rho}_{{\rm red}}|\phi{}_{k}\rangle=\prod_{k}P\left(\phi_{k}\right),\qquad P\left(\phi_{k}\right):=\sqrt{\frac{1}{2\pi\Sigma_{k11}}}\exp\left(-\frac{\phi_{k}^{2}}{2\Sigma_{11}}\right)$ (29)
@@
$\langle\phi|{\rho}_{{\rm red}}|\phi^{\prime}\rangle=\frac{1}{\sqrt{2\pi\Sigma_{11}}}\exp\left[-\frac{1}{2\Sigma_{11}}\phi_{s}^{2}-\frac{\det\Sigma}{2\Sigma_{11}}\phi_{{\rm a}}^{2}+i\frac{\Sigma_{12}}{\Sigma_{11}}\phi_{{\rm s}}\phi_{{\rm a}}\right]$ (30)
@@ 3.3 The second law in cosmology for a Gaussian state
$\frac{\partial}{\partial\eta}\det\Sigma=2{\mathrm{Re}}\,\left[{\cal F}_{k}(\eta,\eta_{0})\right]\Sigma_{11},$ (31)
and (see eqn. (3.37) and the sentence below it in (Burgess:2022nwu, ) ), as pointed out in this reference ${\mathrm{Re}}\,\left[{\cal F}_{k}(\eta,\eta_{0})\right]$is expected to be positive on general grounds [^11] This of course would have followed from the properties of the Lindblad equation, in particular the necessity of preserving the trace and the positivity of the reduced density matrix. However here we do not have strict Markovianity so this is not guaranteed. For a detailed discussion of this including a generalization of the arguments of (Burgess:2022nwu, ), see (Lopez:2025arw, ).. Explicitly in the model that is investigated in detail there,
@@
$\frac{\partial}{\partial\eta}\det\Sigma>0,\qquad{\rm for}\qquad-k\eta\ll 1.$ (32)
So it seems that in this set up $\det\Sigma$ increases without bound for late times and hence the purity appears to go to zero. However as pointed out in the paragraph after eqn. (28) in the previous subsection, purity at the end of inflation for a CMB mode is small but non-zero [^12] Of course as pointed out in footnote 9 since the period of inflation is finite the purity never reaches zero before the end of inflation..
@@
In order to avoid confusion let us stress that the Lindblad equation argument is only used in the first term of (34). In other words the Linblad equation computes $\partial\rho/\partial\eta|_{\Lambda}$ i.e. with a fixed cutoff. The time dependence of the cutoff is taken into account in the second term of (34).
Note added in version 2. After this paper was uploaded to ArXiv, an article Burgess:2025dwm by the first three authors of reference Burgess:2022nwu appeared in which it was argued that the dominant interaction is different from the one considered in Burgess:2022nwu . However while that will affect explicit expressions such as the equation below (31), it must still be linear in the system field as argued in these references. So our discussion which depends only on the general properties of these calculations will not be affected. Indeed the discussion in the next section just depends on there being just one dominant interaction during inflation, a situation that obataines in both these references.
@@ 3.4 Second law in cosmology beyond the Gaussian approximation
$C_{B}^{*}\left(\eta,\eta^{\prime};{\bf x-x^{\prime}}\right)=C_{B}\left(\eta^{\prime},\eta;{\bf x^{\prime}-x}\right)$ (37)
@@
$\frac{\partial\rho}{\partial\eta}\simeq-\frac{1}{2}\int d^{3}x\int d^{3}x^{\prime}G^{2}(\eta)\left(\left\{v\left(\eta,{\bf x}\right)v\left(\eta,{\bf x}^{\prime}\right),\rho(\eta)\right\}-2v\left(\eta,{\bf x}\right)\rho(\eta)v\left(\eta,{\bf x}^{\prime}\right)\right)C_{B}\left({\bf x-x^{\prime}}\right)$ (39)
This has the form of a time-dependent Lindblad equation Martin:2018zbe . From here we can compute the variation of the quantum entropy $S=-{\rm tr}\,\rho\,{\rm ln}\rho$ given by $\frac{dS}{d\eta}=-{\rm tr}\,\frac{d\rho}{d\eta}\ln\rho$.
Now let us assume that the background is spatially compact so that the density matrix can be expanded in a eigenbasis labelled by $\alpha,\beta$ etc. For instance in the absence of interaction (assumed to be the case in the far past) we have $\rho=\prod_{k_{\alpha}>\Lambda}\int_{\phi_{k}}\langle\phi_{k_{\alpha}}|\left\{\phi_{k^{\prime}}\right\}\rangle\langle\{\phi_{k^{\prime}}\}|\phi_{k_{\alpha}}\rangle$, where the $k_{\alpha}$’s take discrete set of values taking space to be compact as before. With time evolution this basis will of course get transformed, but given that the density matrix for the system is still Hermitian there will still be some basis in which $\rho$ will be diagonalized. To simplify the notation we label this basis $\left\{|\alpha\rangle\right\}$ and take it to be discrete with $\alpha|\beta>=\delta_{\alpha\beta}$. So $\rho=\sum_{\alpha}p_{\alpha}|\alpha\rangle\langle\alpha|,$ $\sum_{\alpha}p_{\alpha}=1,\,0<p_{\alpha}<1$.
@@
$\frac{\partial S}{\partial\eta}|_{\Lambda}$ $\simeq\int d^{3}x\int d^{3}x^{\prime}G^{2}\left(\eta\right)\left(\sum_{\alpha\beta}(v(\eta,{\bf x})v(\eta,{\bf x}^{\prime}))_{\alpha\beta}\left(p_{\alpha}\ln\rho_{\alpha}\right)\delta_{\alpha\beta}+p_{\alpha}\left((v(\eta,{\bf x})v(\eta,{\bf x}^{\prime})\right)_{\alpha\beta}\left(\ln p_{\alpha}\right)\delta_{\beta\alpha}\right$.
@@
$=\sigma_{\beta\alpha}$ (40)
@@
$\frac{\partial S}{\partial\eta}|_{\Lambda}$ $\simeq G^{2}\sum_{\alpha\beta}\sigma_{\alpha\beta}\left(p_{\alpha}-p_{\beta}\right)\ln p_{\alpha}=\sum_{\alpha\beta}\sigma_{\beta\alpha}\left(p_{\beta}-p_{\alpha}\right)\ln p_{\beta}$
@@
In the second line above we used (40) and in the last line the positive semi-definiteness of $\sigma$, (41). Also the quantity $\left(p_{\alpha}-p_{\beta}\right)\left(\ln p_{\alpha}-\ln p_{\beta}\right)$ is positive since the log is a monotonic function of its argument. We then conclude that entropy increases monotonically.